# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 2.2. On the hull and the correcting curve
In order to define the hull and to obtain its usesful properties we will need the following notion.
@@
Let $G:=\mathbf{G}(\mathbb{R})$. We say that a closed subgroup $H$ is observable in $G$ if there is a finite-dimensional rational representation $V$ and an $v\in V$ such that $H=\{g\in G:g.v=v\}$.
The notion of observable groups was introduced [Mostow_et_al_observable_groups]. More precisely, [Mostow_et_al_observable_groups] defines observable groups as the algebraic subgroups of algebraic groups for which every finite dimensional rational representation extends to a finite dimensional rational representation of the ambient group. The above definition turns out equivalent to the extension property [Mostow_et_al_observable_groups, Theorem 8] (see [Observable_groups_over_fields, Theorem 9] for this statement for non-algebraically closed fields).
We note:
  • [Observable_group_grosshans, Corollary 2.8]: If $H$ is an algebraic subgroup over $\mathbb{R}$ whose radical is unipotent, then $H$ is observable.
Definition 2.3.
We say that a curve $\varphi\subset G$ is bounded modulo a closed subgroup $H\leq G$ if there exists a compact subset $K\subseteq G$ such that $\varphi\subseteq KH$. Alternatively, $\varphi$ is bounded modulo $H$ if its image in $G/H$ is contained in a compact subset of $G/H$.
Here is our main theorem of this section.
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ is a continuous, non-contracting curve definable in an o-minimal structure $\mathscr{S}$. Then, there exists a unique closed connected (with respect to the real topology) subgroup $H_{\varphi}\leq G$ of the smallest dimension such that $H_{\varphi}$ is generated by one-parameter unipotent subgroup and such that $\varphi$ is bounded modulo $H_{\varphi}$. Moreover, the following properties hold:
@@
    The subgroup $H_{\varphi}$ is minimal in the following sense: if $H^{\prime}$ is an observable group such that $\varphi$ is bounded modulo $H^{\prime}$, then $H_{\varphi}\leq H^{\prime}$. Moreover, if $V$ is a rational representation, and $v\in V$ is such that $\{\varphi(t).v:t\geq 0\}$ is bounded in $V$, then $\varphi$ is bounded modulo the isotropy group of $v$ and $H_{\varphi}.v=\{v\}$.
@@
    If $V$ is a rational finite dimensional representation of $G$ over $\mathbb{R}$, then for $v\in V$ either $H_{\varphi}.v=\{v\}$ or $\lim_{t\to\infty}\beta(t)\varphi(t).v=\infty$.
@@
Let $G=\mathbf{G}(\mathbb{R})$ and and let $\varphi:[0,\infty)\to G$ be a continuous non-contracting curve definable in an o-minimal structure $\mathscr{S}$. We define the hull $H_{\varphi}$ to be the smallest connected closed subgroup generated by unipotent elements such that $\varphi$ is bounded modulo $H_{\varphi}$, and we call a curve $\beta:[0,\infty)\to G$ a correcting curve if $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$.
@@
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a curve whose entry functions $\varphi_{i,j}$ generate an algebra which only consists of either constant functions or functions diverging to $\infty$, e.g., when $t\mapsto\varphi_{i,j}(t)$ are polynomials for all $i,j$. Then it follows by Theorem 2.4 that $\varphi(0)^{-1}\varphi([0,\infty))\subset H_{\varphi}$. A correcting curve is given by the constant curve $\beta(t)=\varphi(0)^{-1},\forall t\geq 0$.
@@
Let $G_{1},G_{2}$ be the real points of algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a rational homomorphism. Suppose that $\varphi:[0,\infty)\to G_{1}$ is a non-contracting curve definable in an o-minimal structure $\mathscr{S}$. Then $f(H_{\varphi})=H_{f\circ\varphi}$.
@@
We start with showing that $H_{f\circ\varphi}\subseteq f(H_{\varphi})$. Let $\beta:[0,\infty)\to G_{1}$ be a correcting curve of $\varphi$. Since $H_{\varphi}$ is generated by unipotents, we get that $f(H_{\varphi})$ is generated by unipotents. Moreover, the curve $f\circ\varphi$ is bounded modulo $f(H_{\varphi})$. In fact, $t\mapsto f(\beta(t))$ is bounded, and for all $t\geq 0$:
@@
We now show the other inclusion. Since $\text{zcl}(H_{f\circ\varphi})$ is observable, we may choose a rational representation $\rho:G_{2}\to\text{GL}(V)$ and a $v\in V$ such that
@@
Since $f\circ\varphi$ is bounded modulo $H_{f\circ\varphi}$, we get that the trajectory
@@
is bounded in $V$. Now, $\eta:=\rho\circ f:G_{1}\to\text{GL}(V)$ is a representation of $G_{1}$ such that $t\mapsto\eta(\varphi(t))v,~{}t\geq 0$ is bounded. By Theorem 2.4(1), we obtain that $\rho(f(H_{\varphi}))v=\eta(H_{\varphi})v=\{v\}$. That is, $f(H_{\varphi})$ is contained in the isotropy group of $v$ which equals to $\text{zcl}(H_{f\circ\varphi})$. ∎
@@
We begin with the following “curve correcting” lemma.
@@
Consider an algebraic action of $G:=\mathbf{G}(\mathbb{R})$ on an affine or projective variety $\mathcal{Z}$. Let $\varphi:[0,\infty)\to G$ be a continuous curve definable in an o-minimal structure $\mathscr{S}$. Suppose that $x\in\mathcal{Z}$ and $g_{0}\in G$ are such that
$\lim_{t\to\infty}\varphi(t).x=g_{0}.x.$
Then there exists a continuous $\mathscr{S}$-definable curve $\delta:[0,\infty)\to G$ such that $\lim_{t\to\infty}\delta(t)=e$, and
$\delta(t)\varphi(t).x=g_{0}.x,\forall t\geq 0.$
@@
$\mathcal{D}$ :=$:=\{(t,g):t>0,g\in G,\|g-I_{n}\|<1\text{ and }g_{0}^{-1}\varphi(t).x=g.x\}.$
@@
$\lim_{t\to\infty}g_{0}^{-1}\varphi(t).x=x,$
we get that for all $t$ large enough, say $t\geq T_{0}$, there exists $g\in G$ such that $(t,g)\in\mathcal{D}$. Using the choice function theorem [Geom_cat_and_o_min_str, Theorem 4.5], there is a definable curve $\tilde{\delta}(t):[T_{0},\infty)\to G$ such that $(t,\tilde{\delta}(t))\in\mathcal{D},~{}\forall t\geq T_{0}$. In particular,
$\tilde{\delta}(t).x=g_{0}^{-1}\varphi(t).x,\forall t\geq T_{0},$
and $\tilde{\delta}$ is bounded. Since the curve $\tilde{\delta}(t)$ is bounded, by the Monotonicity Theorem [Geom_cat_and_o_min_str, Theorem 4.1] we have that $\lim_{t\to\infty}\tilde{\delta}(t)=\delta_{0}$. Note that $\delta_{0}.x=x$. Consider, $\delta(t):=g_{0}(\tilde{\delta}(t)\delta_{0}^{-1})^{-1}g_{0}^{-1}$ for $t\geq T_{0}$. Then,
$\delta(t)\varphi(t).x=g_{0}.x,\forall t\geq T_{0}.$
By cell decomposition [Geom_cat_and_o_min_str, Theorem 4.2], $\delta(t)$ is continuous in $[T^{\prime}_{0},\infty)$. We now extend $\delta$ to the interval $[0,T^{\prime}_{0}]$ by putting
$\delta(t):=\delta(T^{\prime}_{0})\varphi(T^{\prime}_{0})\varphi(t)^{-1},t\leq T^{\prime}_{0}.$
Then $\delta(t)\varphi(t).x=g_{0}.x,\forall t\geq 0$. Since $\varphi$ is continuous and definable, $\delta:[0,\infty)\to G$ is continuous, definable and bounded. ∎
We now observe the following.
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable is an o-minimal structure $\mathscr{S}$. Suppose that $V$ is a finite dimensional rational representation and assume that there is an $v\in V\smallsetminus\{0\}$ such that $\varphi(t).v$ is bounded. Then, there exists a $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t).v=g_{0}.v.$
In particular, there exists a bounded, continuous curve $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ definable in $\mathscr{S}$, such that $\beta(t)\varphi(t).v=v,\forall t\geq 0$.
@@
Since $\varphi:[0,\infty)\to G$ is definable, and since $V$ is a rational representation, it follows that $\varphi(t).v$ is a bounded definable curve. By the monotonicity theorem (see [Geom_cat_and_o_min_str, Section 4]), it follows that there is an $v^{\prime}\in V$ such that $\lim_{t\to\infty}\varphi(t).v=v^{\prime}$. We claim that $v^{\prime}\in G.v$. Denote $S:=\text{zcl}(G.v)\smallsetminus G.v$, and suppose for contradiction that $v^{\prime}\in S$. Then, by Lemma A.1 which is a slight modification of [Kem78, Lemma 1.1(b)], there exists a rational representation $W$ of $G$ and an $G$-equivariant polynomial map $P:V\to W$ such that $P(v)\neq 0$ and $P(S)=0$. But then
$\lim_{t\to\infty}\varphi(t).P(v)=P(\varphi(t).v)=P(v^{\prime})=0,$
which is a contradiction since $\varphi$ is non-contracting. Thus, we conclude that there is an $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t).v=g_{0}.v.$
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in an o-minimal structure. Let $H_{1},H_{2}\leq G$ be two observable groups such that the image of $\varphi$ is bounded in $G/H_{i}$ for $i\in\{1,2\}$. Then, the image of $\varphi$ is bounded in $G/(H_{1}\cap H_{2})$.
@@
For $i\in\{1,2\}$, let $V_{i}$ be a finite dimensional rational representations such that $H_{i}=\{g\in G:g.v_{i}=v_{i}\}$ where $v_{i}\in V_{i}$. Consider the direct-sum representation $V_{1}\oplus V_{2}$, and notice that
$H_{1}\cap H_{2}=\{g\in G:g.(v_{1},v_{2})=(v_{1},v_{2})\}.$
Since the image of $\varphi$ is bounded in $G/H_{i}$, $\varphi(t).v_{i}$ is bounded for all $i\in\{1,2\}$. In particular, $\varphi(t).(v_{1},v_{2})$ is bounded in $V_{1}\oplus V_{2}$. Then, by Lemma 2.10 there is a bounded continuous curve $\delta:[0,\infty)\to G$ such that $\delta(t)\varphi(t).(v_{1},v_{2})=(v_{1},v_{2})$ for all $t\geq 0$. Namely, $\delta(t)\varphi(t)\in H_{1}\cap H_{2}$ for all $t\geq 0$, which shows that claim. ∎
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve definable in an o-minimal structure. Then there exists a unique observable subgroup $\tilde{H}_{\varphi}\leq G$ of the smallest dimension such that $\varphi$ is bounded in $G/\tilde{H}_{\varphi}$, and if $H$ is an observable subgroup such that $\varphi$ is bounded in $G/H$, then $\tilde{H}_{\varphi}\leq H$.
@@
Let $H_{0}$ be a Zariski connected observable group of the smallest dimension such that $\varphi$ is bounded in $G/H_{0}$. If $H$ is any connected observable group such that $\varphi$ is bounded in $G/H$, then by Lemma 2.11, we get that $\varphi$ is bounded in $G/(H_{0}\cap H)$. If $H$ does not include $H_{0}$, then $H\cap H_{0}$ is a proper subgroup of dimension strictly smaller than $\dim H_{0}$, which is impossible. ∎
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in an o-minimal structure. Let $\tilde{H}_{\varphi}\leq G$ be the minimal observable group such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Denote $H_{\varphi}:=\tilde{H}_{\varphi}^{0}$, the identity component of $\tilde{H}_{\varphi}$.
Then, there exists a bounded, continuous curve $\beta:[T_{0},\infty)\to G$ such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$. Moreover, if $V$ is a finite-dimensional rational representation of $G$ and $v\in V$ is such that $\varphi(t).v$ is bounded, then $\beta(t)\varphi(t).v=v,\forall t\geq T_{0}$.
@@
By Lemma 2.10, we obtain a continuous definable curve $\tilde{\beta}:[0,\infty)\to\text{SL}(n,\mathbb{R})$ such that $\tilde{\beta}(t)\varphi(t)\in\tilde{H}_{\varphi}$ for all $t\geq 0$. Since $\varphi$ is continuous, it is contained in a connected component of $H_{\varphi}$. By choosing a suitable $h_{0}\in H_{\varphi}$ we get that for $\beta(t):=h_{0}\tilde{\beta}(t)$ it holds $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$. Now, if $V$ is a finite-dimensional rational representation such that $\varphi(t).v$ is bounded, then by Lemma 2.10, we conclude that $\varphi$ is bounded modulo the observable group $H:=\{g\in G:g.v=v\}$. Hence, by minimality of $\tilde{H}_{\varphi}$, we get that $\beta(t)\varphi(t)\in H_{\varphi}\leq H$. ∎
In order to establish Theorem 2.4 it remains to prove that the identity component of the above minimal observable group is generated by one-parameter unipotent subgroups.
@@
Let $G:=\mathbf{G}(\mathbb{R})$. Suppose that $\varphi:[0,\infty)\to G$ be a non-contracting, continuous curve, definable in an o-minimal structure, and let $\tilde{H}_{\varphi}$ be the smallest observable subgroup such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Then $H_{\varphi}=\tilde{H}_{\varphi}^{0}$ is generated by one-parameter unipotent subgroups.

### v2

@@ 2.2. On the hull and the correcting curve
In what follows, $G:=\mathbf{G}(\mathbb{R})$ is the real points of an algebraic group $\mathbf{G}$ over $\mathbb{R}$. To define the hull and obtain its useful properties, we will need the following notion.
@@
We say that a Zariski closed subgroup $H$ is observable in $G$ if there is a finite-dimensional rational representation $V$ over $\mathbb{R}$ and $v\in V$ such that $H=\{g\in G:g\cdot v=v\}$.
The notion of observable groups was introduced [Mostow_et_al_observable_groups]. More precisely, [Mostow_et_al_observable_groups] defines observable groups as the algebraic subgroups of algebraic groups for which every finite-dimensional rational representation extends to a finite-dimensional rational representation of the ambient group. The above definition turns out to be equivalent to the extension property [Mostow_et_al_observable_groups, Theorem 8] (see [Observable_groups_over_fields, Theorem 9] for this statement for non-algebraically closed fields). The following known result will be useful, see e.g. [Observable_group_grosshans, Corollary 2.8].
Lemma 2.3.
Let $H$ be an algebraic subgroup of an algebraic group $G$ over $\mathbb{R}$. If the radical of $H$ is unipotent, then $H$ is an observable subgroup of $G$.
Remark 2.4.
Suppose that $L$ is a closed subgroup of $\text{SL}(n,\mathbb{R})$ generated by unipotent one-parameter subgroups. Then $L=\text{zcl}(L)^{0}$ and the radical of $\text{zcl}(L)$ is unipotent (see [Shah1991, Lemma 2.9]). Hence, by Lemma 2.3, $\text{zcl}(L)$ is observable in $\text{SL}(n,\mathbb{R})$.
In what follows, we say that a curve $\varphi([0,\infty))\subseteq G$ is bounded modulo a closed subgroup $H\subseteq G$ if there exists a compact subset $K\subseteq G$ such that $\varphi([0,\infty))\subseteq KH$.
We fix an o-minimal structure $\mathscr{S}$ for the rest of the following two subsections. We stress that the assumption of polynomial boundedness on $\mathscr{S}$ is unnecessary here.
Here is our main theorem in this section.
@@
Let $\varphi:[0,\infty)\to G$ be a continuous, non-contracting curve definable in $\mathscr{S}$. Then, there exists a unique closed connected (in the Hausdorff topology) subgroup $H_{\varphi}\subseteq G$ of the smallest dimension such that $H_{\varphi}$ is generated by one-parameter unipotent subgroup and such that $\varphi$ is bounded modulo $H_{\varphi}$. Moreover, the following properties hold:
@@
    The subgroup $H_{\varphi}$ is minimal in the following sense: if $H^{\prime}$ is an observable group such that $\varphi$ is bounded modulo $H^{\prime}$, then $H_{\varphi}\subseteq H^{\prime}$. Moreover, if $V$ is a rational representation, and $v\in V$ is such that $\{\varphi(t)\cdot v:t\geq 0\}$ is bounded in $V$, then $\varphi$ is bounded modulo the isotropy group of $v$ and $H_{\varphi}\cdot v=\{v\}$.
@@
    If $V$ is a rational finite-dimensional representation of $G$ over $\mathbb{R}$, then for any $v\in V$, either $H_{\varphi}\cdot v=\{v\}$ or $\lim_{t\to\infty}\beta(t)\varphi(t)\cdot v=\infty$.
@@
Let $\varphi:[0,\infty)\to G$ be a continuous non-contracting curve definable in $\mathscr{S}$. We define the hull $H_{\varphi}$ to be the smallest connected closed subgroup generated by unipotent elements such that $\varphi$ is bounded modulo $H_{\varphi}$, and we call a curve $\beta:[0,\infty)\to G$ a correcting curve if $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$.
Remark 2.7.
In view of Remark 2.4, $H_{\varphi}=\text{zcl}(H_{\varphi})^{0}$ and $\text{zcl}(H_{\varphi})$ is the observable subgroup of smallest dimension modulo which $\varphi$ is bounded.
@@
Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a definable curve whose entry functions $\varphi_{i,j}$ generate an algebra which consists of functions that are either constant or diverging to $\infty$, e.g., when $t\mapsto\varphi_{i,j}(t)$ are polynomials for all $i,j$. Then, by Theorem 2.5, $\varphi(0)^{-1}\varphi([0,\infty))\subseteq H_{\varphi}$. A correcting curve is given by the constant curve $\beta(t)=\varphi(0)^{-1}$ for all $t\geq 0$.
@@
Let $G_{1}$ and $G_{2}$ be the real points of algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a rational homomorphism. Suppose that $\varphi:[0,\infty)\to G_{1}$ is a non-contracting curve definable in $\mathscr{S}$. Then $f(H_{\varphi})=H_{f\circ\varphi}$.
@@
We start by showing that $H_{f\circ\varphi}\subseteq f(H_{\varphi})$. Let $\beta:[0,\infty)\to G_{1}$ be a correcting curve of $\varphi$. Since $H_{\varphi}$ is generated by unipotents, we get that $f(H_{\varphi})$ is generated by unipotents. Moreover, the curve $f\circ\varphi$ is bounded modulo $f(H_{\varphi})$. In fact, $t\mapsto f(\beta(t))$ is bounded, and for all $t\geq 0$:
@@
We proceed to show the other inclusion. Since $\text{zcl}(H_{f\circ\varphi})$ is observable, we choose a rational representation $\rho:G_{2}\to\text{GL}(V)$ and a $v\in V$ such that
@@
Since $f\circ\varphi$ is bounded modulo $H_{f\circ\varphi}$, the trajectory,
@@
is bounded in $V$. Now, $\eta:=\rho\circ f:G_{1}\to\text{GL}(V)$ is a representation of $G_{1}$ such that $t\mapsto\eta(\varphi(t))v,~{}t\geq 0$ is bounded. By Theorem 2.5(1), we obtain that $\rho(f(H_{\varphi}))v=\eta(H_{\varphi})v=\{v\}$. That is, $f(H_{\varphi})$ is contained in the isotropy group of $v$ which is equal to $\text{zcl}(H_{f\circ\varphi})$. Since $f(H_{\varphi})$ is generated by unipotent one-parameter subgroups,
$f(H_{\varphi})\subseteq\text{zcl}(H_{f\circ\varphi})^{0}=H_{f\circ\varphi}.$
@@
We begin with a “curve correcting” lemma.
@@
Consider an algebraic action of $G$ on the $\mathbb{R}$ points of an affine or projective variety $\mathcal{Z}$ over $\mathbb{R}$. Let $\varphi:[0,\infty)\to G$ be a continuous curve definable in $\mathscr{S}$. Suppose that $x\in\mathcal{Z}$ and $g_{0}\in G$ are such that
$\lim_{t\to\infty}\varphi(t)\cdot x=g_{0}\cdot x.$
Then, there exists a continuous $\mathscr{S}$-definable curve $\delta:[0,\infty)\to G$ such that
$\lim_{t\to\infty}\delta(t)=e\text{ and }\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\ \forall t\geq 0.$
@@
$\mathcal{D}$ :=$:=\{(t,g):t>0,\ g\in G,\|g-I_{n}\|<1\text{ and }g_{0}^{-1}\varphi(t)\cdot x=g\cdot x\}.$
@@
$\lim_{t\to\infty}g_{0}^{-1}\varphi(t)\cdot x=x,$
we get that for all $t$ large enough, say $t\geq T_{0}$, there exists $g\in G$ such that $(t,g)\in\mathcal{D}$. Using the choice function theorem, see Theorem B.7, there is a definable curve $\tilde{\delta}(t):[T_{0},\infty)\to G$ such that $(t,\tilde{\delta}(t))\in\mathcal{D}$ for all $t\geq T_{0}$. In particular,
$\tilde{\delta}(t)\cdot x=g_{0}^{-1}\varphi(t)\cdot x,\,\forall t\geq T_{0},$
and $\tilde{\delta}$ is bounded. Since the curve $\tilde{\delta}(t)$ is bounded, by the monotonicity theorem, see Theorem B.5, we have that $\lim_{t\to\infty}\tilde{\delta}(t)=\delta_{0}$. Note that $\delta_{0}\cdot x=x$. Let $\delta(t):=g_{0}(\tilde{\delta}(t)\delta_{0}^{-1})^{-1}g_{0}^{-1}$ for all $t\geq T_{0}$. Then,
$\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\forall t\geq T_{0}.$
By the monotonicity theorem, $\delta(t)$ is continuous in $[T^{\prime}_{0},\infty)$, for some $T^{\prime}_{0}\geq T_{0}$. We now extend $\delta$ to the interval $[0,T^{\prime}_{0}]$ by putting
$\delta(t):=\delta(T^{\prime}_{0})\varphi(T^{\prime}_{0})\varphi(t)^{-1},\,\forall t\leq T^{\prime}_{0}.$
Then $\delta(t)\varphi(t)\cdot x=g_{0}\cdot x,\forall t\geq 0$. Since $\varphi$ is continuous and definable, $\delta:[0,\infty)\to G$ is continuous, definable and bounded. ∎
@@
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Suppose that $V$ is a finite-dimensional rational representation of $G$ and $v\in V\setminus\{0\}$ is such that $\varphi(t)\cdot v$ is bounded. Then, there exists $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t)\cdot v=g_{0}\cdot v.$
In particular, there exists a bounded, continuous curve $\beta:[0,\infty)\to G$ definable in $\mathscr{S}$, such that $\beta(t)\varphi(t)\cdot v=v$ for all $t\geq 0$.
@@
Since $\varphi:[0,\infty)\to G$ is definable, and since $V$ is a rational representation, $\varphi(t)\cdot v$ is a bounded definable curve. By the monotonicity theorem (see Theorem B.5), $\lim_{t\to\infty}\varphi(t)\cdot v=v^{\prime}\in V$. We claim that $v^{\prime}\in G\cdot v$. Denote $S:=\text{zcl}(G\cdot v)\setminus G\cdot v$, and suppose for contradiction that $v^{\prime}\in S$. Then, by Lemma A.1, which is a slight modification of [Kem78, Lemma 1.1(b)], there exists a rational representation $W$ of $G$ and $G$-equivariant polynomial map $P:V\to W$ such that $P(v)\neq 0$ and $P(S)=0$. In particular, $P(v^{\prime})=0$. Then
$\lim_{t\to\infty}\varphi(t)\cdot P(v)=P(\lim_{t\to\infty}\varphi(t)\cdot v)=P(v^{\prime})=0,$
which is a contradiction, as $\varphi$ is non-contracting. Thus, we conclude that there exists $g_{0}\in G$ such that
$\lim_{t\to\infty}\varphi(t)\cdot v=g_{0}\cdot v.$
@@
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Let $H_{1}$ and $H_{2}$ be observable groups of $G$ such that the image of $\varphi$ is bounded in $G/H_{i}$ for $i\in\{1,2\}$. Then, the image of $\varphi$ is bounded in $G/(H_{1}\cap H_{2})$.
@@
For $i\in\{1,2\}$, let $V_{i}$ be a finite-dimensional rational representation such that $H_{i}=\{g\in G:g\cdot v_{i}=v_{i}\}$ where $v_{i}\in V_{i}$. Consider the direct-sum representation $V_{1}\oplus V_{2}$, and notice that
$H_{1}\cap H_{2}=\{g\in G:g\cdot(v_{1},v_{2})=(v_{1},v_{2})\}.$
Since the image of $\varphi$ is bounded in $G/H_{i}$, $\varphi(t)\cdot v_{i}$ is bounded for all $i\in\{1,2\}$. In particular, $\varphi(t)\cdot(v_{1},v_{2})$ is bounded in $V_{1}\oplus V_{2}$. Then, by Lemma 2.12, there is a bounded continuous curve $\delta:[0,\infty)\to G$ such that $\delta(t)\varphi(t)\cdot(v_{1},v_{2})=(v_{1},v_{2})$ for all $t\geq 0$. Namely, $\delta(t)\varphi(t)\in H_{1}\cap H_{2}$ for all $t\geq 0$, and therefore the image of $\varphi$ is bounded in $G/(H_{1}\cap H_{2})$. ∎
@@
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve definable in $\mathscr{S}$. Then there exists a unique observable subgroup $\tilde{H}_{\varphi}\subseteq G$ of minimal dimension such that $\varphi$ is bounded in $G/\tilde{H}_{\varphi}$. Moreover, if $H$ is an observable subgroup such that $\varphi$ is bounded in $G/H$, then $\tilde{H}_{\varphi}\subseteq H$.
@@
Let $H_{0}$ be a Zariski connected observable group of the smallest dimension such that $\varphi$ is bounded in $G/H_{0}$. If $H$ is any connected observable group such that $\varphi$ is bounded in $G/H$, then by Lemma 2.13, we get that $\varphi$ is bounded in $G/(H_{0}\cap H)$. If $H_{0}\not\subseteq H$, then $H\cap H_{0}$ is a proper subgroup of dimension strictly smaller than $\dim H_{0}$, which leads to a contradiction. ∎
@@
Suppose that $\varphi:[0,\infty)\to G$ is a non-contracting, continuous curve, definable in $\mathscr{S}$. Let $\tilde{H}_{\varphi}\subseteq G$ be the minimal observable group such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Denote $H_{\varphi}:=\tilde{H}_{\varphi}^{0}$, the identity component of $\tilde{H}_{\varphi}$.
Then, there exists a bounded, continuous definable curve $\beta:[T_{0},\infty)\to G$ such that $\beta(t)\varphi(t)\in H_{\varphi}$ for all $t\geq 0$. Moreover, if $V$ is a finite-dimensional rational representation of $G$ and $v\in V$ is such that $\varphi(t)\cdot v$ is bounded, then $\beta(t)\varphi(t)\cdot v=v$ for all $t\geq T_{0}$.
@@
By Lemma 2.12, we obtain a continuous definable curve $\tilde{\beta}:[0,\infty)\to G$ such that $\tilde{\beta}(t)\varphi(t)\in\tilde{H}_{\varphi}$ for all $t\geq 0$. Since $\varphi$ is continuous, it is contained in a connected component of $\tilde{H}_{\varphi}$. By choosing a suitable $h_{0}\in\tilde{H}_{\varphi}$ we get that for $\beta(t):=h_{0}\tilde{\beta}(t)$ it holds $\beta(t)\varphi(t)\in\tilde{H}^{0}_{\varphi}=H_{\varphi}$, for all $t\geq 0$. Now, if $V$ is a finite-dimensional rational representation such that $\varphi(t)\cdot v$ is bounded, then by Lemma 2.12, we conclude that $\varphi$ is bounded modulo the observable group $H:=\{g\in G:g\cdot v=v\}$. Hence, by minimality of $\tilde{H}_{\varphi}$, we get that $\beta(t)\varphi(t)\in H_{\varphi}\subseteq H$. ∎
The last property in Theorem 2.5 remaining to prove is that the identity component of the observable group of smallest dimension is generated by one-parameter unipotent subgroups. This is given by the following.
@@
Let $\varphi:[0,\infty)\to G$ be a $\mathscr{S}$-definable, non-contracting, continuous curve, and let $\tilde{H}_{\varphi}$ be the smallest observable subgroup such that $\varphi$ is bounded modulo $\tilde{H}_{\varphi}$. Then $H_{\varphi}=\tilde{H}_{\varphi}^{0}$ is generated by one-parameter unipotent subgroups.